A gas expands from 4.55 L to 6.7 L mL against an external pressure of 1.52 atm. In addition, 1254 J is emitted as heat. What is the change in energy of the burning of the fuel?
Answer #1

LA) Change in energy, AU=? 00 = 9+w 9 = -1254 ] ( energy released) WE - PSV E-P(V2-Y,) = -1,52 atin (6.7 - 4.55) + = -1.52 x